A bullet can go through a block of 1 centimetre by what factor should it be increased so that it goes through four such blocks
1
Expert's answer
2020-08-06T16:46:22-0400

If we assume that kinetic energy of the bullet fully converts into energy "E_b" required to go trough a block, for one block the situation will look like


"\frac{mv^2}{2}=E_b."

For 4 blocks, the equation is


"\frac{mu^2}{2}=4E_b,\\\space\\
u=\sqrt{\frac{8E_b}{m}}=\sqrt{\frac{8(mv^2/2)}{m}}=2v."

To go through 4 blocks, the speed can be increased by a factor of 2.